# Decoding-Time Language Model Alignment with Multiple Objectives

### v1

@@ 5.2 Results
Reddit Summary. By supervised fine-tuning a Llama2-7B model on Summarize-from-Feedback dataset nisan2020summarize , we obtain the reference policy $\pi_{\operatorname{ref}}$. And then we obtain $\pi_{1},\pi_{2}$ by tuning $\pi_{\operatorname{ref}}$ using PPO for two off-shelf reward models which measures summary quality and faithfulness, respectively. Then we show Pareto frontiers of MOD, RS, and MORLHF in Figure 3, with preference weightings $w\in\left\{(i/10,1-i/10):i\in\{0,1,\ldots,10\}\right\}$, demonstrating the superiority of MOD over baselines.
Helpful Assistant. By supervised fine-tuning a Llama2-7B model on Anthropic-HH dataset Bai2022TrainingAH , we obtain the reference policy $\pi_{\operatorname{ref}}$. And then we obtain $\pi_{1},\pi_{2},\pi_{3}$ by tuning $\pi_{\operatorname{ref}}$ using PPO for three off-shelf reward models which evaluate helpfulness, harmlessness, and humor, respectively. The Pareto frontiers of MOD, RS and MORLHF for each two-objective pairs are shown in Figure 3. MOD lags behind MORLHF in a certain task, while MORLHF is more costly. We explore the $3$-objective setting on the Helpful Assistant task, demonstrating that MOD can effectively balance advantages of each model and outperforms RS. More results are provided in Appendix F.2.
Figure 4: Safety Alignment. Figures from left to right illustrate $f$-DPO models w.r.t. Reverse KL-divergence, JSD, $0.3$-divergence and $0.5$-divergence, respectively. MODPO is only applicable to KL-divergence, and we report its mean of $3$ seeds. The frontier of MOD lies over RS across all these divergences. Although MODPO exhibits a larger area, MOD is smoother.
Safety Alignment. Based on results reported in fDPO , we mainly focus on $f$-DPO with Reverse KL-divergence, JSD, $0.3$-divergence and $0.5$-divergence in experiments. We deploy an off-shelf Alpaca-7B model as $\pi_{\operatorname{ref}}$ and train $\pi_{1f},\pi_{2f}$ using $f$-DPO on two pair-comparison BeaverTails-$10$beavertails datasets: one is Better and the other is Safer. We show Pareto frontiers of MOD, RS, and MODPO for each $f$-divergence in Figure 4. Experimental results demonstrate that MOD generally outperforms RS across these $f$-divergences. The retraining baseline MODPO is only applicable to Reverse KL-divergence, and MOD is much more steerable and convenient compared with MODPO despite a slight performance gap.
@@
Table 4: Results of MOD combining CodeTülu-2-7B, Tülu-2-HH-13B, and Tülu-2-Ultra-13B, achieving precise control over general capabilities, including safety (Toxigen), coding (Codex), and reasoning ($\ast$ COT). MOD with $w=(0.75,0.1,0.15)$ reduces Toxigen to nearly 0% and achieves 7.9–33.3% improvement across the other three metrics, compared with CodeTülu-2-7B.
@@
figurePerformance of combining three Tülu models. Our combinations (in orange and blue) exhibit better overall performance than single models. [Uncaptioned image]
@@ 6.1 Failures of parameter-merging paradigm
The optimality of the parameter-merging paradigm rewardedsoups ; personalizedsoup primarily relies on reduced reward mis-specification (see 1). The following theorem demonstrates that this hypothesis hardly holds for almost all $f$-divergence regularized policies. See detailed proof in Appendix D.5. {restatable}[]theoremthmrewardsoup For any $f$-divergence satisfying one of the following conditions: (i) $f$ is not a barrier function; (ii) $I_{f}$ is Reverse KL-divergence; (iii) $f$ is a strong-barrier function, with finite roots of
@@
there $\exists N,M\in\mathbb{N}$, $\mathcal{Y}=\{y_{i}\}_{i=1}^{N}$, $\beta\in\mathbb{R}_{+}$, a neural network $nn=\operatorname{softmax}(h_{\theta}(z_{0}))$ where $z_{0}\in\mathbb{R}^{n}$ and $h_{\theta}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{N}$ is a continuous mapping, preference weightings $w\in\Delta^{M-1}$, reference policy $\pi_{\operatorname{ref}}$, and the objectives $J_{1},J_{2},\ldots,J_{M}$ representing reward functions $\mathcal{R}_{1},\mathcal{R}_{2},\ldots,\mathcal{R}_{M}$ w.r.t. $\beta\cdot I_{f}(\cdot\|\pi_{\operatorname{ref}})$, s.t. 1 does not hold.
@@
It is commonly adopted in previous studies ziegler2019finetune ; nisan2020summarize that the network receives the same inputs $z_{0}$. Despite the competitive results exhibited in prior works modelsoup ; rewardedsoups ; personalizedsoup , this theorem reveals that parameter-merging lacks a theoretical guarantee in practical scenarios. Besides, although 1 may hold, the mapping from preference weightings $w$ to the optimal merging weightings $\lambda$ are intricate, and thus simply picking $\lambda$ as $w$ in RS rewardedsoups , can yield sub-optimal results.
Another perspective of the same initialization. We can also look into scenarios where only the parameters of the last several layers of $\pi_{1},\pi_{2},\ldots,\pi_{M}$ can be different from $\pi_{\operatorname{ref}}$. 1) If the last layer is linear projection, then it is equivalent to MOD w.r.t. $\operatorname{KL}\left(\cdot\|\pi_{\operatorname{ref}}\right)$, namely linearly combining the logits. 2) If the last layer is self-attention transformer , then it can be easily hacked by reversing the sign of $Q,K$ matrices in this layer, which does not influence the value of $Q^{\top}K$, but significantly harms the effect of parameter-merging. A motivating example is shown in Appendix F.1.
@@ 6.2 Necessity of barrier function
Extending the results of fDPO to the multi-objective setting, we prove the necessity of $f$ being barrier functions to find an optimal policy $\pi^{\star}$ for multi-objective alignment. See detailed proof in Appendix D.2. {restatable}[]theoremthmwontwork If $f$ is not a barrier function, then for $\forall C\in\mathbb{R}_{+}$, $N\in\mathbb{Z}_{\geq 4}$, $M\in\mathbb{Z}_{\geq 2}$, $\mathcal{Y}=\{y_{i}\}_{i=1}^{N}$, any multi-objective decoding or merging algorithm $\mathcal{A}:\mathcal{S}^{M+1}\times\Delta^{M-1}\rightarrow\mathcal{S}$, there exists a reference policy $\pi_{\operatorname{ref}}$, policies $\{\mathcal{\pi}_{i}\}_{i=1}^{M}$ and $\pi^{\prime}$, reward functions $\{\mathcal{R}_{i}\}_{i=1}^{M}$, preference weightings $w\in\Delta^{M-1}$ and $\beta\in\mathbb{R}_{+}$, s.t. $\pi_{i}$ is the optimal policy for $\mathcal{R}_{i}$ w.r.t. $\beta\cdot I_{f}(\cdot\|\pi_{\operatorname{ref}})$ (see 1), $\forall i\in[M]$, but
@@
Crucial role of the barrier function. We can apply this theorem to any algorithm which solely utilizes base policies, including RS and MOD. And thus, a barrier function regularization is crucial in multi-objective alignment to bridge different policies, though it is intended to prevent degeneration (see Table 3 in DPO ) in single-objective alignment. Additionally, the same as a general barrier in interior point methods convexoptimization , it obviates the need for introducing slack variables as in fDPO . This explains why we should not use non-barrier $f$-divergences such as total variation and chi-squared.
@@ 6.4 Beyond $f$-divergence regularized alignment and multi-objective decoding
While our main results are based on $f$-divergence regularized aligned LMs and aimed at multi-objective decoding, our framework is also applicable to using SFT models and explaining the effectiveness of other existing decoding algorithms. For example, proxy-tuning proxytuning tunes only a smaller LM, then applies the difference between the logits of the small tuned and untuned LMs to shift the predictions of a larger untuned model. Its theoretical justification can be reduced to our framework, under certain assumptions. We provide insights on this line of work proxytuning ; Jailbreak ; huang2024offset and derivations of some other related works liu2024decoding ; modpo in Appendix C.3, further demonstrating the potential for universally applying our approach.
@@ 7 Conclusion
It is also important to acknowledge the limitations of our work. 1) The analysis is primarily based on tabular parametrization, not taking function approximation error into consideration. 2) Decoding from a response-level probability distribution at the token level may lead to degraded performance, which is likely to be alleviated by energy-based approaches cold ; Kumar2021ControlledTG ; zhao2024probabilistic .
@@ Appendix A Impact statement
Appendix B Notations
Here we introduce a set of notations to be used throughout. For any differentiable function $f$, let $\nabla f$ denote its gradient. For any $N\in\mathbb{N}$, we denote the index set $\{1,\cdots,N\}$ as $[N]$. Let $e_{s}$ be the $s_{\text{th}}$ standard basis vector. For any $i,j\in\mathbb{Z}_{\geq 0}$, $\delta_{ij}$ represents the Kronecker delta function [15], which output $1$ if $i=j$ otherwise 00. For any $n\in\mathbb{N}$, $\Delta^{n}$ represents the $n$-dimensional probability simplex $\{(p_{1},\ldots,p_{n+1}):p_{i}\geq 0,\;\forall i\in[n+1],\;\sum_{j=1}^{n+1}p_{j}=1\}$, and $\Delta(X)$ represents the set of probability distributions over a set $X$. $\mathcal{X}$ denotes the prompt set, $\Sigma$ denotes the alphabet set, $\mathcal{Y}\subset\Sigma^{\ast}$ denotes the response set, and the policy set $\mathcal{S}$ is defined as all mappings from $\mathcal{X}$ to $\Delta(\mathcal{Y})$.
@@ C.2 Divergence measures and closed-form policies
We acknowledge that commonly used $f$-divergence measures have been introduced in [46] and show them here for completeness:
@@ C.3 Extended variants
Proxy-tuning [27] & jail-breaking [57]. Based on the claim above, for another base model $\mathcal{M}$, we thus have
@@
which reflects the tuned version of model $\mathcal{M}$. And this is exactly the proxy-tuning approach, validated by extensive experiments in [27]. Reversing the position of $\mathbb{P}_{\mathcal{M}^{+}}$ and $\mathbb{P}_{\mathcal{M}_{-}}$ yields jail-breaking [57]. $\delta$-unlearning [19] is the same.
Multi-objective proxy-tuning. Moreover, it is worth noting that, our method can be applied as a lightweight approach for large-scale models, as a multi-objective extension of proxy-tuning [27]. In particular, to tune a large pre-trained model $\mathcal{M}$, we can first tune $\mathcal{M}^{+}_{1},\mathcal{M}^{+}_{2},\ldots,\mathcal{M}^{+}_{M}$ from a relatively smaller model $\mathcal{M}^{-}$ by PPO, DPO or SFT, and decode $y_{t}$ at timestep $t$ as
@@
DeRa [29]. Given $\mathbb{P}_{\mathcal{M}^{+}}(y|x)\propto\mathbb{P}_{M^{-}}(y|x)\cdot\exp(\frac{1}{\beta}r(y|x))$, then
@@
MODPO [60]. Assuming $\pi_{i}$ is the optimal policy for $\mathcal{R}_{i}$ w.r.t. $\beta\operatorname{KL}\left(\cdot\|\pi_{\operatorname{ref}}\right)$, $\forall i\in[M]$, then the optimal policy for $\sum_{i=1}^{M}w_{i}\mathcal{R}_{i}$ w.r.t. $\beta\operatorname{KL}\left(\cdot\|\pi_{\operatorname{ref}}\right)$, $\pi^{\star}\propto\prod\pi_{i}^{w_{i}}$, is the minimizer of

### v3

@@ 4.2 Results
Reddit Summary. By supervised finetuning a Llama2-7B model on Summarize-from-Feedback dataset nisan2020summarize , we obtain the reference policy $\pi_{\operatorname{ref}}$. And then we obtain $\pi_{1},\pi_{2}$ by tuning $\pi_{\operatorname{ref}}$ using PPO for two off-shelf reward models (see details in Appendix E) which measures summary quality and faithfulness, respectively. Then we show Pareto frontiers of MOD, RS, and MORLHF in Figure 3, with preference weightings $w\in\left\{(i/10,1-i/10):i\in\{0,1,\ldots,10\}\right\}$, demonstrating the superiority of MOD over baselines.
Helpful Assistant. By supervised finetuning a Llama2-7B model on Anthropic-HH dataset Bai2022TrainingAH , we obtain the reference policy $\pi_{\operatorname{ref}}$. And then we obtain $\pi_{1},\pi_{2},\pi_{3}$ by tuning $\pi_{\operatorname{ref}}$ using PPO for three off-shelf reward models (see details in Appendix E) which evaluate helpfulness, harmlessness and humor, respectively. The Pareto frontiers of MOD, RS and MORLHF for each two-objective pairs are shown in Figure 3. MOD prominently beats RS for each reward pair, and lags behind MORLHF in balancing harmlessness and humor, while MORLHF is more costly. We explore the $3$-objective setting on the Helpful Assistant task, demonstrating that MOD can effectively balance advantages of each model and outperforms RS. More results are provided in Appendix F.2.
Figure 4: Safety Alignment. Figures from left to right illustrate $f$-DPO models w.r.t. Reverse KL-divergence, JSD, $0.3$-divergence and $0.5$-divergence, respectively. MODPO is only applicable to KL-divergence, and we report its mean of $3$ seeds. The frontier of MOD generally lies over RS.
Safety Alignment. Based on results reported in fDPO , we mainly focus on $f$-DPO with Reverse KL-divergence, JSD, $0.3$-divergence and $0.5$-divergence in experiments. We deploy an off-shelf Alpaca-7B model as $\pi_{\operatorname{ref}}$ and train $\pi_{1f},\pi_{2f}$ using $f$-DPO on two pair-comparison BeaverTails-$10$beavertails datasets: one is Better and the other is Safer. We show Pareto frontiers of MOD, RS, and MODPO for each $f$-divergence in Figure 4. Experimental results demonstrate that MOD generally outperforms RS across multiple $f$-divergences. The retraining baseline MODPO is only applicable to Reverse KL-divergence, and MOD is much more convenient despite a slight performance gap.
@@
Table 4: Results of MOD combining CodeTülu-2-7B, Tülu-2-HH-13B, and Tülu-2-Ultra-13B, achieving precise control over general capabilities, including safety (Toxigen), coding (Codex), and reasoning ($\ast$ COT). MOD with $w=(0.75,0.1,0.15)$ reduces Toxigen to nearly 0 and achieves 7.9–33.3% improvement across the other three metrics, compared with CodeTülu-2-7B.
@@
tablePerformance of combining three Tülu models. Our combinations (in orange and blue) exhibit better overall performance than single models. [Uncaptioned image]
@@
Notably, for any finite number of objectives, there exists infinite possible weightings. In this experiment, we discretize the weightings space using small grid size like $0.1,0.15,0.3$. Based on this, we randomly set $3$ three combinations without careful picking. Intuitively, the weightings should reflect the users’ preferences on the general objectives that those models are good at.
@@ 5.1 Failures of parameter-merging paradigm
The optimality of the parameter-merging paradigm rewardedsoups ; personalizedsoup primarily relies on reduced reward mis-specification hypothesis (see 1 in Appendix D.1 for definition). The following theorem demonstrates that this hypothesis does not hold for almost all $f$-divergence regularized policies. See detailed proof in Appendix D.5. {restatable}[]theoremthmrewardsoup For any $f$-divergence satisfying one of the following conditions: (i) $f$ is not a barrier function; (ii) $I_{f}$ is Reverse KL-divergence; (iii) $f$ is a strong-barrier function, with finite roots of
@@
$\exists N,M\in\mathbb{N}$, $\mathcal{Y}=\{y_{i}\}_{i=1}^{N}$, $\beta\in\mathbb{R}_{+}$, a neural network $nn=\operatorname{softmax}(h_{\theta}(z_{0}))$ where $z_{0}\in\mathbb{R}^{n}$ and $h_{\theta}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{N}$ is a continuous mapping, preference weightings $w\in\Delta^{M-1}$, reference policy $\pi_{\operatorname{ref}}$, and the objectives $J_{1},J_{2},\ldots,J_{M}$ representing reward functions $\mathcal{R}_{1},\mathcal{R}_{2},\ldots,\mathcal{R}_{M}$ w.r.t. $\beta\cdot I_{f}(\cdot\|\pi_{\operatorname{ref}})$, s.t. 1 does not hold.
@@
It is commonly adopted in previous studies ziegler2019finetune ; nisan2020summarize that the network receives the same inputs $z_{0}$. Despite the competitive results exhibited in prior works modelsoup ; rewardedsoups ; personalizedsoup , this theorem reveals that parameter-merging lacks a theoretical guarantee in practical scenarios. Besides, although 1 may hold, the mapping from preference weightings $w$ to the optimal merging weightings $\lambda$ are intricate, and thus simply picking $\lambda$ as $w$ rewardedsoups , can yield sub-optimal results.
Another perspective of the same initialization. We can also look into scenarios where only the parameters of the last several layers of $\pi_{1},\pi_{2},\ldots,\pi_{M}$ can be different from $\pi_{\operatorname{ref}}$. 1) If the last layer is a linear projection, then it is equivalent to MOD w.r.t. $\operatorname{KL}\left(\cdot\|\pi_{\operatorname{ref}}\right)$, namely linearly combining the logits. 2) If the last layer is self-attention transformer , then it can be easily hacked by reversing the sign of $Q,K$ matrices in this layer, which does not influence the value of $Q^{\top}K$, but significantly harms the effect of parameter-merging. A motivating example is shown in Appendix F.1.
@@ 5.2 Necessity of barrier function
Extending the results of fDPO to the multi-objective setting, we prove the necessity of $f$ being barrier functions to find an optimal policy $\pi^{\star}$ for multi-objective alignment. See detailed proof in Appendix D.2. {restatable}[]theoremthmwontwork If $f$ is not a barrier function, then for $\forall C\in\mathbb{R}_{+}$, $N\in\mathbb{Z}_{\geq 4}$, $M\in\mathbb{Z}_{\geq 2}$, $\mathcal{Y}=\{y_{i}\}_{i=1}^{N}$, any multi-objective decoding or merging algorithm $\mathcal{A}:\mathcal{S}^{M+1}\times\Delta^{M-1}\rightarrow\mathcal{S}$, there exists a reference policy $\pi_{\operatorname{ref}}$, policies $\{\mathcal{\pi}_{i}\}_{i=1}^{M}$ and $\pi^{\prime}$, reward functions $\{\mathcal{R}_{i}\}_{i=1}^{M}$, preference weightings $w\in\Delta^{M-1}$ and $\beta\in\mathbb{R}_{+}$, s.t. $\pi_{i}$ is the optimal policy for $\mathcal{R}_{i}$ w.r.t. $\beta\cdot I_{f}(\cdot\|\pi_{\operatorname{ref}})$ (see 1 in Appendix D.1), $\forall i\in[M]$, but
@@
Crucial role of the barrier function. We can apply this theorem to any algorithm which solely utilizes base policies, including RS and MOD. And thus, a barrier function regularization is crucial in multi-objective alignment to bridge different policies, though it was originally intended to prevent degeneration (see Table 3 in DPO ) in single-objective alignment. Additionally, the same as a general barrier in interior point methods convexoptimization , it obviates the need for introducing slack variables as in fDPO . This explains why we should not use non-barrier $f$-divergences such as total variation and chi-squared.
@@ 5.4 Beyond $f$-divergence regularized alignment and multi-objective decoding
While our main results are based on $f$-divergence regularized aligned LMs and aimed at multi-objective decoding, our framework is also applicable to using SFT models and explaining the effectiveness of other existing decoding algorithms. For example, proxy-tuning proxytuning tunes only a smaller LM, then applies the difference between the logits of the small tuned and untuned LMs to shift the predictions of a larger untuned model. Its theoretical justification is provided by our framework, under certain assumptions. We provide insights on this line of work proxytuning ; Jailbreak ; huang2024offset and derivations of some other related works liu2024decoding ; modpo in Appendix C.3, further demonstrating the potential for universally applying our approach.
6 Related Work
Algorithms for aligning LMs to human preferences. The widely used RLHF (PPO) approach InstructGPT ; nisan2020summarize ; wu2023fine optimizes over rewards with Reverse KL-divergence as a penalty, where the reward models are learned from human preference datasets. DPO DPO leverages the Bradley-Terry assumption BTmodel to directly optimize the same objective on preferences, in a supervised manner. $\Psi$-PO IPO further modifies the reward term to be optimized as other mappings from preference pairs; f-DPO fDPO replaces Reverse KL-divergence with other divergence measures. In addition, there are other efforts exploring alternative objectives and frameworks: SLiC-HF SLiC ; SLiCHF refer to the alignment process as sequence likelihood calibration; SPIN SPIN iteratively improves the model by leveraging synthetically generated data, thereby circumventing the need for human feedback; OPO alignonthefly employs established norms as constraints, achieving training-free alignment; and Lyu et  al. Lyu2024KeepingLA highlight the crucial role of prompt templates. In this work, we mainly focus on RLHF (PPO), DPO, and their extensions.
Decoding-time algorithms for controllable generation. Response-level decoding algorithms sample a whole output $y$ from an anticipated probability distribution $p$. To achieve this goal, energy-based methods are adopted in many works cold ; gradientbasedsampling , which involves continuous optimization for LMs to obtain gradient information. Kumar et  al. Kumar2021ControlledTG view this task as maximizing $\log p(y)$ while satisfying some constraints, and use simultaneous gradient descent to solve it. Token-level decoding algorithms decode token $y_{t}$ at timestep $t$, and are usually more efficient. Among them, Mudgal et  al. controlleddecoding , Liu et  al. Liu2023DontTA deploy value models to guide the decoding process; DeRa liu2024decoding works on hyperparameter re-alignment and proposes the potential of a special case of MOD, while introducing a per-token distribution approximation; proxy-tuning proxytuning ; Jailbreak ; huang2024offset tunes a small model and applies it to steer a larger base model by operating on logits.
Multi-objective LMs alignment. Multi-objective alignment is the task of aligning language models to multiple objectives simultaneously. This is important for managing tradeoffs among different dimensions Vamplew2017HumanalignedAI ; Bai2022TrainingAH and catering to the diverse needs of users personalizedsoup ; dong-etal-2023-steerlm . Approaches for multi-objective alignment fall into the following categories: 1) Retraining. The most natural approach to solve multi-objective alignment is to retrain for a linearly combined multiple reward functions (MORLHF wu2023fine ; Bai2022TrainingAH ). And MODPO modpo retrains the model in a reward-model-free way, by learning a flexible reward representation and directly training on a fixed preference dataset. 2) Parameter-merging. This line of work rewardedsoups ; personalizedsoup ; Lin2023MitigatingTA , represented by rewarded soups (RS), aims at providing a training-free solution which obtains weights of the policy as a linear combination of weights of trained policies for each single objective, inspired by modelsoup and its other applications WARM ; lawson2023merging . Jiang et  al. Jiang2023LLMBlenderEL achieve another kind of model-merging through reranking and fusion on outputs. 3) Preference-conditioned prompting. The preference-conditioned learning approaches Zhu2023ScalingPD ; Basaklar2022PDMORLPM train a policy conditioned on preference weightings to maximize the expected rewards. This concept is reflected in LMs alignment as preference-conditioned prompting: this line of work RiC ; haoxiang2024arithmetic ; guo2024controllable directly presents the preference weightings in prompts after a finetuning process. The latter two paradigms are more efficient, while relying heavily on either the reduced mis-specification hypothesis rewardedsoups or unguaranteed OOD generalization ability Zhou2023FreeFB , posing challenges in terms of interpretability and robustness.
@@ 7 Conclusion
It is also important to acknowledge the limitations of our work. 1) The analysis is primarily based on tabular setting xu2020preference , not taking function approximation error into consideration. 2) Decoding from a response-level probability distribution at the token level may lead to degraded performance, which is likely to be alleviated by energy-based approaches cold ; Kumar2021ControlledTG ; zhao2024probabilistic .
Acknowledgement
SSD acknowledges the support of NSF IIS 2110170, NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, and NSF IIS 2229881. NAS acknowledges the support of NSF IIS 2113530. The authors also thank Yizhong Wang for useful discussions.
@@ Appendix A Impact Statement
Appendix B Notation
Here we introduce a set of notation to be used throughout. For any differentiable function $f$, let $\nabla f$ denote its gradient. For any $N\in\mathbb{N}$, we denote the index set $\{1,\cdots,N\}$ as $[N]$. Let $e_{s}$ be the $s$th standard basis vector. For any $i,j\in\mathbb{Z}_{\geq 0}$, $\delta_{ij}$ represents the Kronecker delta function linearalgebratextbook , which output $1$ if $i=j$ otherwise 00. For any $n\in\mathbb{N}$, $\Delta^{n}$ represents the $n$-dimensional probability simplex $\{(p_{1},\ldots,p_{n+1}):p_{i}\geq 0,\;\forall i\in[n+1],\;\sum_{j=1}^{n+1}p_{j}=1\}$, and $\Delta(X)$ represents the set of probability distributions over a set $X$. $\mathcal{X}$ denotes the prompt set, $\Sigma$ denotes the alphabet set, $\mathcal{Y}\subset\Sigma^{\ast}$ denotes the response set, and the policy set $\mathcal{S}$ is defined as all mappings from $\mathcal{X}$ to $\Delta(\mathcal{Y})$.
@@ C.2 Divergence measures and closed-form policies
We acknowledge that commonly used $f$-divergence measures have been introduced in fDPO and show them here for completeness:
@@ C.3 Extended variants
Proxy-tuning proxytuning & jail-breaking Jailbreak . Based on the claim above, for another base model $\mathcal{M}$, we thus have
@@
which reflects the tuned version of model $\mathcal{M}$. And this is exactly the proxy-tuning approach, validated by extensive experiments in proxytuning . Reversing the position of $\mathbb{P}_{\mathcal{M}^{+}}$ and $\mathbb{P}_{\mathcal{M}_{-}}$ yields jail-breaking Jailbreak . $\delta$-unlearning huang2024offset is the same.
Multi-objective proxy-tuning. Moreover, it is worth noting that, our method can be applied as a lightweight approach for large-scale models, as a multi-objective extension of proxy-tuning proxytuning . In particular, to tune a large pre-trained model $\mathcal{M}$, we can first tune $\mathcal{M}^{+}_{1},\mathcal{M}^{+}_{2},\ldots,\mathcal{M}^{+}_{M}$ from a relatively smaller model $\mathcal{M}^{-}$ by PPO, DPO or SFT, and decode $y_{t}$ at timestep $t$ as
@@
DeRa liu2024decoding . Given $\mathbb{P}_{\mathcal{M}^{+}}(y|x)\propto\mathbb{P}_{M^{-}}(y|x)\cdot\exp(\frac{1}{\beta}r(y|x))$, then
@@
MODPO modpo . Assuming $\pi_{i}$ is the optimal policy for $\mathcal{R}_{i}$ w.r.t. $\beta\operatorname{KL}\left(\cdot\|\pi_{\operatorname{ref}}\right)$, $\forall i\in[M]$, then the optimal policy for $\sum_{i=1}^{M}w_{i}\mathcal{R}_{i}$ w.r.t. $\beta\operatorname{KL}\left(\cdot\|\pi_{\operatorname{ref}}\right)$, $\pi^{\star}\propto\prod\pi_{i}^{w_{i}}$, is the minimizer of